OpenStudy (anonymous):

6.A local grocery store stacks the soup cans in such a way that each row has 2 fewer cans than the row below it. If there are 32 cans on the bottom row, how many total cans are on the bottom 14 rows? (1 point). This is an arithmetic series because?

OpenStudy (anonymous):

don't you mean how many cans are on the top row,? any way, to figure this out we will have to write an equation, it would be something, arithmetically like row, can, So if we start with 32 cans, then C=(R*2)-32, This means that if you want to find the 14th row, then simply multiply 14 * 2 and subtract it by 32. Sorry it took me long.

OpenStudy (anonymous):

well the directions say i dont have to solve it it pretty much just wants me to tell why its an arithmetic series

OpenStudy (anonymous):

I just showed you, C=(R*2)-32 is arithmetic because for every row, the number of cans decreases.

OpenStudy (anonymous):

row0=32 row1= 32 -2= row0 -2 row2 = row1-2 in genearl row_n = row_{n-1} -2 And that is what an arithmetic series is?

OpenStudy (anonymous):

ooo lol

OpenStudy (anonymous):

thankxx